OpenStudy (anonymous):

Suppose that ABCD is a parallelogram, and A=(-1,2), B=<2,b> C=(5,-1), D= What are the coordinates of a and b?

OpenStudy (amoodarya):

hint : A+C=B+D it means (XA,YA)+(XC,YC)=(XB,YB)+(XD,YD) put the values in relation soon you will find the answer